OpenStudy (anonymous):

Find h when v=679.84 and r=6 using the formula V=π r2h

OpenStudy (campbell_st):

so you have \[679.84 = \pi \times 6^2 \times h... or... 679.84 = 36\pi h\] just make h the subject.